OpenStudy (anonymous):

find the distance between (4/5, -1/2) and (2,0). express answer in simplest form.

OpenStudy (amistre64):

the distance is the same as applying the pythag thrm

OpenStudy (amistre64):

(distance of x)^2 + (distance of y)^2 = (distance between the points)^2

OpenStudy (anonymous):

so distance = root((2-0.8)^2 + (0.4)^2))

OpenStudy (amistre64):

( x , y ) ( 2 , 0 ) -(4/5, -1/2) ----------- ( 6/5 , 1/2 ) \[\left(\frac{6}{5}\right)^2+\left(\frac{1}{2}\right)^2=distance^2\]

OpenStudy (amistre64):

\[\frac{36}{25}+\frac{1}{4}=d^2\] \[\frac{144+25}{100}=d^2\] \[\sqrt{\frac{169}{100}}=d\] \[\frac{13}{10}=d=1.3\] if I kept track of it all

OpenStudy (anonymous):

so its 1.3?

OpenStudy (anonymous):

yes..sorry

OpenStudy (amistre64):

as far as I can see, I dont see any mistakes; so im going with 1.3